The average public elementary school has 468 students with a standard deviation of 87.If a random sample of 38 public elementary schools is selected,what is the probability that the number of students enrolled is between 445 and 485

1
Expert's answer
2022-06-12T14:58:26-0400

"P(445<\bar X<485)=P(\frac{445-468}{\frac{87}{\sqrt{38}}}<Z<\frac{485-468}{\frac{87}{\sqrt{38}}})="

"=P(-1.63<Z<1.20)=P(Z<<1.20)-P(Z<-1.63)=0.8333."